find the point of intersection.
y= x-4
y= -9x+6

Answer by rfer(16322)

x-4=-9x+6
x+9x=6+4
10x=10
x=1
---------------
y=-9*1+6
y=-9+6
y=-3
--------------
(1,-3)
-----------
-3=1-4